### v1

@@ 1 Introduction
Artificial Intelligence (AI) and Machine Learning have revolutionized the connectivity and autonomy of vehicles, including trucks. With the integration of sensors, cameras, and sophisticated onboard systems, machine learning algorithms can analyze vast amounts of real-time data, allowing trucks to make intelligent decisions on the road. The continuous advancements in machine learning continue to push the boundaries of connectivity and autonomy, transforming the trucking industry towards a more efficient and intelligent future.
@@
A widely used machine learning framework in the context of autonomous systems is Reinforcement Learning (RL). RL overcomes the challenges of traditional search based methods such as A* search which lack the ability to generalize to unknown situations in a non-deterministic environment and are computationally expensive [9]. RL has been increasingly used for solving complex problems related to autonomous driving, such as navigation, trajectory planning, collision avoidance, and behavior planning [10, 11, 12]. The application of this framework to autonomous trucks is a relatively new area of research. An important contribution in this direction is the work [13], which implements an RL framework for autonomous truck driving in SUMO. They study how a Bayesian RL technique, based on an ensemble of neural networks with additional randomized prior functions (RPF), can be used to estimate the uncertainty of decisions in autonomous driving. On the other hand, more results exist when we consider autonomous driving of passenger cars. The study in [14] develops a cooperative adaptive cruise control (CACC) using RL for securing longitudinal following of a front vehicle using vehicle-to-vehicle communication. The CACC system has a coordination layer that is responsible for the selection of high-level actions, e.g., lane changing or secure vehicle following. The action layer, which is a policy-gradient RL agent, must achieve this action by selecting the appropriate low-level actions that correspond to the vehicle’s steering, brakes, and throttle. Another study in [15] proposes a Supervised Actor-Critic (SAC) [16, 17] approach for optimal control of ACC. Their framework contains an upper level controller based on SAC, which chooses desired acceleration based on relative velocity and distance parameters of leading and following vehicles. A low level controller receives the acceleration signal and transfers it to the corresponding brake and/or throttle control action. In [18], the authors develop a Deep Reinforcement Learning (DRL) framework for ACC and compare the results with Model Predictive Control. In all these studies, the RL agent performs step control of acceleration or other driving maneuvers such as braking and steering. For this reason, it can take a longer time for the agent to achieve maximum speed even though there is no leading vehicle in front, making the process apparently inefficient.
@@
There is only limited prior work where the RL agent is trained to learn high level actions such as choosing the safe gap with the leading vehicle and the actual speed control is performed by a low level controller. The work in [19] develops a similar system that aims to achieve simultaneous optimization of traffic efficiency, driving safety, and driving comfort through dynamic adaptation of the inter-vehicle gap. It suggests a dual RL agent approach - the first RL agent is designed to find and adapt the optimal time-to-collision (TTC) threshold based on rich traffic information, including both macroscopic and microscopic traffic data obtained from the surrounding environment and the second RL agent is designed to derive an optimal inter-vehicle gap that maximizes traffic flow, driving safety, and comfort at the same time. In [20], authors develop a Deep Deterministic Policy Gradient (DDPG) - Proportional Integral Derivative (PID) controller for longitudinal control of vehicle platooning. They use the DDPG algorithm to optimize the $K_{p}$, $K_{d}$ and $K_{i}$ constants for the PID controller which controls the desired speed. These works mainly focus on longitudinal control of the vehicles in a platoon, and the effects of lane change maneuvers are not studied sufficiently.
In this study, we develop ACC, together with lane change maneuvers for an autonomous truck in a highway scenario. We first develop a baseline architecture for tactical decision making and control based on the previous work [13, 21]. Here, the speed control and lane change decision making are done by an RL agent. With the baseline architecture, we show how adding relevant features to the observation space improves the performance of the agent. Then, we train a refined architecture where we separate the high-level decision making and low-level control actions between the RL agent and low-level controllers based on physical models. Compared to the existing architectures in the literature, our RL agent only needs to learn a policy to choose high level actions such as the gap to be maintained. Interestingly, we find this optimized action space improves the learning compared to speed control actions regardless of the RL algorithm used. We evaluate the performance with three different RL algorithms: Deep Q-Network (DQN) [22], Advantage Actor-Critic (A2C) [23] and Proximal Policy Optimization (PPO) [24].
We design two distinct reward functions to train the RL agents. The first one focuses on safety aspects, while the second, more complex one, is based on Total Cost of Operation (TCOP). This TCOP-oriented reward function is realistic, encompassing actual expenses like energy consumption and human resources. TCOP serves as a comprehensive metric covering various operational expenses in autonomous trucking. By introducing a TCOP-based reward function, we aim to optimize driving strategies for sustainable and economically viable operations, in line with our broader objective of fostering cost-effective and environmentally sustainable transportation solutions. Furthermore, our refined RL architecture, combined with this TCOP-centric reward function, establishes a robust framework for future research endeavours, aimed at enhancing the economic viability and operational efficiency of autonomous driving systems.
This work is an extension of our previous work [25] wherein we additionally, (i) investigate training of RL agent with a complex TCOP based reward function consisting of realistic costs and revenue values with and without normalization (ii) extend our RL framework by incorporating curriculum learning techniques and compare the results with the non-curriculum learning approach. In Curriculum Reinforcement Learning (CRL) [26, 27], the RL agent starts learning simpler tasks and is gradually exposed to complex tasks. Effectiveness of curriculum learning has been investigated in a number of problems related to autonomous driving, though different from our setting. The study in [28] introduces multi-stage learning in complex driving environments. Different stages include a diverse set of starting locations, varying weather conditions, and dense traffic scenarios. The action space consists of accelerator or brake values and steering angles. The reward function penalizes collisions, following the wrong route, and exceeding speed limits, guiding the agents towards safe and efficient driving behaviors. Despite not yet reaching state-of-the-art performance, the study demonstrates consistent driving behavior across different scenarios and robustness to weather changes. The papers [29] and [30] utilize a curriculum learning approach for overtaking in autonomous driving. [29] employs a two-stage successive learning progression, where agents first learn to drive as fast as possible and then master the skill of overtaking efficiently. On the other hand, [30] introduces a three-stage curriculum learning process where agents first learn to race, then focus on overtaking maneuvers, and finally refine their skills to race at high speeds while avoiding collisions. Both of these works demonstrate that curriculum learning boosts convergence and contributes to a better final policy.
@@
All our results are obtained using Simulation of Urban Mobility (SUMO) [31], which is a widely used simulation platform for the conceptual development of autonomous vehicles. The tailored RL environment based on SUMO that we developed for heavy vehicle combination driving in highway traffic is provided as open access, which offers a great open source framework for investigating various RL methods in complex settings. [^1] Source Code: https://github.com/deepthi-pathare/Autonomous-truck-sumo-gym-env
@@ 2.2.1 MDP Formulation
  • Reward Function(R): A basic reward function focused on safety aspects is designed to motivate the agent to drive at maximum speed and reach the target without any hazardous situations. This reward function is similar to that in [13] except that a reward for reaching the target is added, whereas the penalty for emergency braking is not considered for simplicity. The reward at each time step is given by (1).
@@
    Here $v_{t}$ is the velocity of the vehicle at time step $t$ and $max\_v$ is the maximum velocity of the vehicle. $I$ is an indicator function, which takes a value of 1 when the corresponding condition is satisfied and $P$ and $R$ are the corresponding penalty and reward values respectively. The possible conditions are lane change ($l$), collision ($c$), near collision ($nc$), driving outside the road ($o$) and reaching the target ($tar$). $T$ is the total time it takes to reach the target. The parameter values used are given in Table 1.
@@ 2.2.2 Reinforcement Learning Algorithms
    $L^{CLIP}(\theta)=$ (6)
    $\hat{{E}}_{t}\Big{[}\min\Big{(}r_{t}(\theta)\hat{A}_{t},\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}\Big{)}\Big{]}$
@@ 2.4 Total Cost of Operation
$P_{l}$
0.1
$P_{c}$
1000 euros
$P_{nc}$
1000 euros
$P_{o}$
1000 euros
$R_{tar}$
2.78 euros
$C_{el}$
0.5 euro per kwh
$C_{dr}$
50 euro per hour
$\Delta t$
1 s
$m$
40000 kg
$C_{d}$
0.36
$A_{f}$
10 $m^{2}$
$\rho_{\text{air}}$
1.225 $kg/m^{3}$
$g$
9.81 $m/s^{2}$
$C_{r}$
0.005
In the reward function (8), the penalties are defined as the average cost incurred during hazardous situations. The average cost considered here is the own risk payment which is the portion of an insurance claim made by the vehicle owner or the insured for any loss and or damage that occurs when submitting a claim. Similarly, the reward $R_{tar}$ denotes the revenue that can be achieved by the truck when it completes the target. Revenue is computed as the total expected cost with $20\%$ profit in an ideal scenario where the truck drives with an average speed of $22m/s$ and zero acceleration. The total length of travel is 2200 meters and hence the time to reach the target will be $100s$. Based on (9) and parameter values from Table 2, the total energy consumed can be computed as $1.85$ kwh. Then, the total energy cost will be $1.85$ kwh ×0.5=0.925$\times$ 0.5=0.925× 0.5 = 0.925 euros. The total driver cost for $100s$ will be $1.39$ euros and the total cost becomes $0.925+1.39=2.315$ euros. Further, adding the $20\%$ profit gives the net revenue $2.78$ euros, which is used in the reward function along with a weight $W_{tar}$.
@@ 2.5 Curriculum Reinforcement Learning
Curriculum learning is a training strategy where the model is gradually exposed to increasingly complex tasks or difficult examples over time. This approach is based on the intuition that starting with simpler tasks allows the model to build foundational knowledge and gradually progress to more challenging tasks, which can help improve learning efficiency. We combine curriculum learning with RL in the new system architecture and investigate its effectiveness in training the RL agent with the complex reward function based on TCOP. Here, the curriculum for training the RL agent is designed as follows:
@@
In this CRL approach, we update the reward function in each curriculum whereas the RL environment, state and action space remain the same. In this work, we have redefined TCOP based reward function as shown in Equation 11 by removing the penalty for lane change and the weight component $W_{tar}$ such that the reward function only contains actual cost and revenue values during a truck’s operation. Training an agent with realistic reward functions can lead to policies that are more aligned with human expectations.
@@ 3.1 Performance improvement based on selection of states
Reached target successfully
61%
70.6%
Driven successfully, but not reached target within maximum steps
0%
0%
Terminated by collision or driving outside road
39%
29.4%
Average speed
18.17 m/s
19.43 m/s
Average distance travelled
1503.93 m
1667.86 m
Average executed steps
88.43
89.65
@@ 3.2 Performance comparison of the two architectures
Reached target successfully
70.6%
97.8%
Driven successfully, but not reached the target within maximum steps
0%
0.6%
Terminated by collision or driving outside the road
29.4%
1.6%
Average speed
19.43 m/s
18.56 m/s
Average distance travelled
1667.86 m
2178.37 m
Average executed steps
89.65
128.75
@@ 3.3.1 Comparison with different weight values
In this analysis, we apply the TCOP-based reward function in Equation 8 to the new architecture in order to assess whether the agent can successfully acquire a driving strategy that is both safe and cost-effective. The episodic reward is evaluated with varying weights $W_{tar}$ assigned to the target completion reward $R_{tar}$, as depicted in Fig. 7. As mentioned in Section II, in an ideal scenario, the cost would amount to 2.314 euros, while the revenue ($R_{tar}$) would be 2.78 euros. Consequently, in this ideal case, the episodic reward would equal $2.78-2.315=0.46$ when $W_{tar}$ is 1. This value is relatively small, making it challenging for the agent to learn how to reach the target, as illustrated in Fig. 7. Moreover, Table 5 demonstrates that the agent exhibits slow driving behavior and struggles to reach the target. However, it becomes evident that increasing the weight assists the agent in learning to drive at higher speeds and successfully reach the target. Notably, as the weight is increased, the total cost of operation—comprising energy cost and driver cost—is minimized to 3.58 euros. This cost is lower than the cost incurred when the ego vehicle was entirely controlled by SUMO, using the default Krauss car following model and LC2013 lane change model in the same traffic environment, which resulted in a cost of 3.85 euros.
Figure 7: Learning of TCOP-based rewards in the new architecture with different $W_{tar}$ values using PPO.
Table 5: Evaluation of new architecture with TCOP based reward function, having different weights, $W_{tar}$.
Evaluation Metric
$W_{tar}=1$
$W_{tar}=10$
$W_{tar}=36$
Reached target successfully
3.6%
1%
98.8%
Driven successfully, but not reached the target within maximum steps
93.4%
96.8%
0.4%
Terminated by collision or driving outside road
3%
2.2%
0.8%
Average speed
2.39 m/s
1.56 m/s
18.24 m/s
Average distance travelled
861.64 m
580.65 m
2187.62 m
Average executed steps
477.02
486.11
126.46
Average energy cost
0.26 euros
0.16 euros
1.82 euros
Average driver cost
6.62 euros
6.75 euros
1.76 euros
Average tcop
6.88 euros
6.91 euros
3.58 euros

### v2

@@ 1 Introduction
Artificial Intelligence (AI) and Machine Learning have revolutionized the connectivity and autonomy of vehicles, including trucks. With the integration of sensors, cameras, and sophisticated onboard systems, machine learning algorithms can analyze vast amounts of real-time data, allowing trucks to make intelligent decisions on the road. Continuous advancements in machine learning continue to push the boundaries of connectivity and autonomy, transforming the trucking industry towards a more efficient and intelligent future.
@@
A widely used machine learning framework in the context of autonomous systems is Reinforcement Learning (RL). RL overcomes the challenges of traditional search based methods such as A* search, which lack the ability to generalize to unknown situations in a non-deterministic environment and are computationally expensive [9]. RL methods also have significant advantages over traditional car-following and lane change models which are based on predefined rules and assumptions. Behavior of traditional models are deterministic and do not adapt beyond their programmed logic. On the otherhand, advanced RL methods such as Deep Reinforcement Learning (DRL) can adapt to non-deterministic environments with varying surrounding vehicle behaviors and make informed decisions in complex scenarios by processing high-dimensional state representations. RL has been increasingly used to solve complex problems related to autonomous car driving, such as navigation, trajectory planning, collision avoidance, and behavior planning [10, 11, 12]. The application of this framework to autonomous trucks is a relatively new area of research. An important contribution in this direction is the work [13], which implements an RL framework for autonomous truck driving in SUMO. They study how a Bayesian RL technique, based on an ensemble of neural networks with additional randomized prior functions (RPF), can be used to estimate the uncertainty of decisions in autonomous driving. On the other hand, more results exist when we consider autonomous driving of passenger cars. The study in [14] develops a cooperative adaptive cruise control (CACC) using RL for securing longitudinal following of a front vehicle using vehicle-to-vehicle communication. The CACC system has a coordination layer that is responsible for the selection of high-level actions, e.g., lane changing or secure vehicle following. The action layer, which is a policy-gradient RL agent, must achieve this action by selecting the appropriate low-level actions that correspond to the vehicle’s steering, brakes, and throttle. Another study in [15] proposes a Supervised Actor-Critic (SAC) [16, 17] approach for optimal control of ACC. Their framework contains an upper level controller based on SAC, which chooses desired acceleration based on relative velocity and distance parameters of leading and following vehicles. A low level controller receives the acceleration signal and transfers it to the corresponding brake and/or throttle control action. In [18], the authors develop a DRL framework for ACC and compare the results with Model Predictive Control. In all these studies, the RL agent performs step control of acceleration or other driving maneuvers such as braking and steering. For this reason, it can take longer time for the agent to achieve maximum speed even though there is no leading vehicle in front, making the process apparently inefficient.
@@
There is also only limited prior work available where the RL agent is trained to learn high level actions such as choosing the safe gap with the leading vehicle and the actual speed control is performed by a low level controller. The work in [19] develops a similar system that aims to achieve simultaneous optimization of traffic efficiency, driving safety, and driving comfort through dynamic adaptation of the inter-vehicle gap. It suggests a dual RL agent approach - the first RL agent is designed to find and adapt the optimal time-to-collision (TTC) threshold based on rich traffic information, including both macroscopic and microscopic traffic data obtained from the surrounding environment and the second RL agent is designed to derive an optimal inter-vehicle gap that maximizes traffic flow, driving safety, and comfort at the same time. In [20], authors develop a Deep Deterministic Policy Gradient (DDPG) - Proportional Integral Derivative (PID) controller for longitudinal control of vehicle platooning. They use the DDPG algorithm to optimize the $K_{p}$, $K_{d}$ and $K_{i}$ constants for the PID controller which controls the desired speed. Both these studies mainly focus on longitudinal control of vehicles, and the effects of lane change maneuvers are not studied.
We develop ACC, together with lane change maneuvers for an autonomous truck in a highway scenario. We propose an architecture that combines RL with low-level controllers and separate the high level and low level decision making between them to improve safety and efficiency. The results of this architecture are compared to a baseline architecture that solely relies on RL to perform actions. We evaluate the performance with three different RL algorithms: Deep Q-Network (DQN) [21], Advantage Actor-Critic (A2C) [22] and Proximal Policy Optimization (PPO) [23]. We design two reward functions for training the RL agents. The first focuses on safety, while the second is based on Total Cost of Operation (TCOP), covering realistic expenses like energy consumption and human resources. We investigate different training methods to handle the complex and realistic reward function based on TCOP. This approach establishes a strong foundation for future research into improving the economic viability and operational efficiency of autonomous driving systems.
@@
This work is an extension of our previous work [24] wherein we additionally, (i) investigate training of RL agent with the complex TCOP based reward function consisting of realistic costs and revenue values with and without normalization (ii) extend our RL framework by incorporating curriculum learning techniques [25, 26] and compare the results with the non-curriculum learning approach. All our results are obtained using Simulation of Urban Mobility (SUMO) [27], which is a widely used simulation platform for the conceptual development of autonomous vehicles. The tailored RL environment based on SUMO that we developed for heavy vehicle combination driving in highway traffic is provided as open access, which offers a great open source framework for investigating various RL methods in complex settings. [^1] Source Code: https://github.com/deepthi-pathare/Autonomous-truck-sumo-gym-env
@@ 2.2.1 MDP Formulation
  • Reward Function(R): We design a basic reward function focused on safety aspects to motivate the agent to drive at maximum speed and reach the target without any hazardous situations. This reward function is similar to that in [13] except that a reward for reaching the target is added, whereas the penalty for emergency braking is not considered for simplicity. The reward at each time step is given by (1).
@@
    Here $v_{t}$ is the velocity of the vehicle at time step $t$ and max_vmax\_v is the maximum velocity of the vehicle. $I$ is an indicator function, which takes a value of 1 when the corresponding condition is satisfied and $P$ and $R$ are the corresponding penalty and reward values respectively. The possible conditions are lane change ($l$), collision ($c$), near collision ($nc$), driving outside the road ($o$) and reaching the target (tartar). $T$ is the total time it takes to reach the target. The parameter values used are given in Table 1.
@@ 2.2.2 Reinforcement Learning Algorithms
    $L^{CLIP}(\theta)=\hat{{E}}_{t}\Big[\min\Big(r_{t}(\theta)\hat{A}_{t},\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}\Big)\Big]$ (6)
@@ 2.4 Total Cost of Operation
$P_{l}$ 0.1
$P_{c}$ 1000 euros
$P_{nc}$ 1000 euros
$P_{o}$ 1000 euros
$R_{tar}$ 2.78 euros
$C_{el}$ 0.5 euro per kwh
$C_{dr}$ 50 euro per hour
$\Delta t$ 1 s
$m$ 40000 kg
$C_{d}$ 0.36
$A_{f}$ 10 $m^{2}$
$\rho_{\text{air}}$ 1.225 $kg/m^{3}$
$g$ 9.81 $m/s^{2}$
$C_{r}$ 0.005
In the reward function (8), the penalties are defined as the average cost incurred during hazardous situations. The average cost considered here is the own risk payment which is the portion of an insurance claim made by the vehicle owner or the insured for any loss and or damage that occurs when submitting a claim. Similarly, the reward $R_{tar}$ denotes the revenue that can be achieved by the truck when it completes the target. Revenue is computed as the total expected cost with $20\%$ profit in an ideal scenario where the truck drives with an average speed of $22m/s$ and zero acceleration. The total length of travel is 2200 meters and hence the time to reach the target will be $100s$. Based on (9) and parameter values from Table 2, the total energy consumed can be computed as $1.85$ kwh. Then, the total energy cost will be $1.85$ kwh $\times 0.5=0.925$ euros. The total driver cost for $100s$ will be $1.39$ euros and the total cost becomes $0.925+1.39=2.315$ euros. Further, adding the $20\%$ profit gives the net revenue $2.78$ euros, which is used in the reward function along with a weight $W_{tar}$. We also added weight components $W_{c}$, $W_{nc}$, $W_{o}$ for the penalties for collision, near collision and driving outside the road condition respectively.
@@ 2.5 Curriculum Reinforcement Learning
Curriculum Reinforcement Learning (CRL) is a training strategy where the model is gradually exposed to increasingly complex tasks or difficult examples over time. This approach is based on the intuition that starting with simpler tasks allows the model to build foundational knowledge and gradually progress to more challenging tasks, which can help improve learning efficiency.
Effectiveness of curriculum learning has been investigated in a number of problems with complex tasks related to autonomous driving, though different from our setting. The study in [33] introduces multi-stage learning in complex driving environments. Different stages include a diverse set of starting locations, varying weather conditions, and dense traffic scenarios. The action space consists of accelerator or brake values and steering angles. The reward function penalizes collisions, following the wrong route, and exceeding speed limits, guiding the agents towards safe and efficient driving behaviors. The papers [34] and [35] utilize a curriculum learning approach for overtaking in autonomous driving. [34] employs a two-stage successive learning progression, where agents first learn to drive as fast as possible and then master the skill of overtaking efficiently. On the other hand, [35] introduces a three-stage curriculum learning process where agents first learn to race, then focus on overtaking maneuvers, and finally refine their skills to race at high speeds while avoiding collisions. Both of these works demonstrate that curriculum learning boosts convergence and contributes to a better final policy.
The above mentioned studies do not address the problem of strategic decision making for optimizing cost, efficiency and safety at the same time. We address this by combining curriculum learning with RL in the new system architecture and investigate its effectiveness in training the RL agent with the complex reward function based on TCOP. Here, the curriculum for training the RL agent is designed as follows:
@@
In this CRL approach, we update the reward function in each curriculum whereas the RL environment, state and action space remain the same. In this work, we have redefined TCOP based reward function as shown in Equation 11 by removing the penalty for lane change and the weight components such that the reward function only contains actual cost and revenue values during a truck’s operation. Training an agent with realistic reward functions can lead to policies that are more aligned with human expectations.
@@ 3.1 Performance improvement based on selection of states
Reached target successfully 61% 70.6%
Driven successfully, but not reached target within maximum steps 0% 0%
Terminated by collision or driving outside road 39% 29.4%
Average speed 18.17 m/s 19.43 m/s
Average distance travelled 1503.93 m 1667.86 m
Average executed steps 88.43 89.65
@@ 3.2 Performance comparison of the two architectures
Reached target successfully 70.6% 97.8%
Driven successfully, but not reached the target within maximum steps 0% 0.6%
Terminated by collision or driving outside the road 29.4% 1.6%
Average speed 19.43 m/s 18.56 m/s
Average distance travelled 1667.86 m 2178.37 m
Average executed steps 89.65 128.75
@@ 3.3.1 Comparison with different weight values
In this analysis, we apply the TCOP-based reward function in Equation 8 to the new architecture in order to assess whether the agent can successfully acquire a driving strategy that is both safe and cost-effective. The episodic reward is evaluated with varying weights $W_{tar}$ assigned to the target completion reward $R_{tar}$, as depicted in Fig. 7. As mentioned in Section II, in an ideal scenario, the cost would amount to 2.315 euros, while the revenue ($R_{tar}$) would be 2.78 euros. Consequently, in this ideal case, the episodic reward would equal 2.782.315=0.462.78-2.315=0.46 when $W_{tar}$ is 1. This value is relatively small, making it challenging for the agent to learn how to reach the target. Furthermore, the penalties for hazardous situations such as collision, near collision and driving outside the road are 1000 euros which is comparatively very high. This leads to a risk for the agent to focus only on avoiding such situations and do not care about reaching the target, for example by driving with very low speed. This issue can be overcome by carefully choosing the weight values for the penalty and revenue components. We observed good results by keeping the weight value for penalties $W_{c}$, $W_{nc}$, $W_{o}$ = 0.1 and increasing the weight value for revenue $W_{tar}$ to 20 as illustrated in Figure 7. From Table 5, it becomes evident that increasing the weight $W_{tar}$ assists the agent in learning to drive at higher speeds and successfully reach the target. Notably, as the weight $W_{tar}$ is increased, the total cost of operation—comprising energy cost and driver cost—is minimized to 3.74 euros. This cost is lower than the cost incurred when the ego vehicle was entirely controlled by SUMO, using the default Krauss car following model and LC2013 lane change model in the same traffic environment, which resulted in a cost of 3.85 euros.
Figure 7: Learning of TCOP-based rewards in the new architecture with PPO using different $W_{tar}$ values and $W_{c}$, $W_{nc}$, $W_{o}$ = 0.1.
Table 5: Evaluation of new architecture with TCOP based reward function, using different $W_{tar}$ values and $W_{c}$, $W_{nc}$, $W_{o}$ = 0.1
Evaluation Metric $W_{tar}=1$ $W_{tar}=5$ $W_{tar}=10$ $W_{tar}=20$
Reached target successfully 2.2% 13.6% 68.2% 99.2%
Driven successfully, but not reached the target within maximum steps 95.2% 83.6% 29.6% 0%
Terminated by collision or driving outside road 2.6% 2.8% 2.2% 0.8%
Average speed 1.75 m/s 2.78 m/s 11.16 m/s 18.36 m/s
Average distance travelled 615.9 m 848.2 m 1864.7 m 2197.4 m
Average executed steps 482 455 265 122
Average energy cost 0.17 euros 0.28 euros 1.1 euros 2.05 euros
Average driver cost 6.7 euros 6.32 euros 3.68 euros 1.69 euros
Average tcop 6.87 euros 6.6 euros 4.78 euros 3.74 euros